OpenStudy (anonymous):

How does 36/4 + 75/4 = 6/2? It gave me the answer, but I dont understand how it = 6/2..? Help please!!! XC

OpenStudy (anonymous):

That's not possible :o If youre talking about adding simple fractions, that's not true

OpenStudy (anonymous):

well is there any other way to get that result?

OpenStudy (anonymous):

this is only a small portion of the problem

OpenStudy (texaschic101):

36/4 + 75/4 = 111/4 = 27 3/4

OpenStudy (anonymous):

well theres a few ways you can get your answer. first by adding the numerators then the denominators then simplifying. so, 36+75=111 and 4+4=8 you cant possible get 6/2 with that

OpenStudy (texaschic101):

your not multiplying, your adding...denominators stay the same

OpenStudy (anonymous):

I got the same answer @texaschic101

OpenStudy (anonymous):

ah yes. sorry. but still in order to go from 4 to 2 you have to divide by 2. and if that's the case 111/2 is 55.5 not 6

OpenStudy (texaschic101):

there is no other way to get the answer....6/2, which equals 3, is not the answer

OpenStudy (texaschic101):

Kaggler...if you got 111/4 and turned it into a mixed number, and got 27 3/4, then that is the correct answer

OpenStudy (anonymous):

I'm just trying to figure out how they got 6/2.. maybe it was a typing error?

OpenStudy (texaschic101):

are there any parenthesis in the problem ?

OpenStudy (anonymous):

nope. this is for solving vertex form.. do you know about that? i solved the rest of the problem correctly but this part of it, i can't match it up to 6/2

OpenStudy (texaschic101):

sorry...I do not know about that

OpenStudy (anonymous):

its okay, the other parts of the problem doesnt have anything to do with this anyways lol. i think there was just a typing error, thank you!